OpenStudy (anonymous):

In the neighborhood park there is a circular grassy picnic area, but half of the grass has died. The city wants to replant the grass, but needs to know the square footage of the damaged area. If the circular picnic area is 150 feet across, what is the square footage of the damaged area? Let π = 3.14.

OpenStudy (anonymous):

IS IT 75 ????

OpenStudy (anonymous):

a) 8831.25 b)177662.5 c)35325 d) 70650

OpenStudy (mathmate):

"circular picnic area is 150 feet across" means the diameter is 150', or the radius is 75'. Area of whole picnic site = pi r^2 Area of damaged site = (1/2)pi r^2 =(1/2)(22/7)*75^2 =8839 sq.ft